OpenStudy (herpderp):

ok ok I have a real question. how come when you divide fractions, you have to flip them and junk?

OpenStudy (apples):

\[ \frac{a}{\frac{b}{c}} = a \cdot (\frac{b}{c})^{-1} = a \cdot (b \cdot c^{-1})^{-1} = a \cdot b^{-1} \cdot c = ac \cdot b^{-1} = \frac{ac}{b}\]

OpenStudy (apples):

Sorry, it cut off. the last part is \[= \frac{ac}{b}\]

OpenStudy (herpderp):

OOOOOOOOOHHHHH

OpenStudy (radar):

dividing by 2 is the same as multiplying by 1/2. Try it and see.

OpenStudy (herpderp):

wow you're so right what a great answer way better than apples